carlosesgirl:

how do you find slope with these points (1,5) and (8, - -7)

Timmyspu:

You need a equation to put these points in to solve for the slope. Do you know what that could be?

carlosesgirl:

(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu You need a equation to put these points in to solve for the slope. Do you know what that could be? \(\color{#0cbb34}{\text{End of Quote}}\) no I don't

Timmyspu:

Well then do you know how to put those points on your graph and see if they have a slope?

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu Well then do you know how to put those points on your graph and see if they have a slope? \(\color{#0cbb34}{\text{End of Quote}}\) no

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carlosesgirl \(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu Well then do you know how to put those points on your graph and see if they have a slope? \(\color{#0cbb34}{\text{End of Quote}}\) no \(\color{#0cbb34}{\text{End of Quote}}\)

carlosesgirl:

here's the problems

BraydenLove:

Use Desmos.com

xxDeppressionxx:

What grade is you in?

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx What grade is you in? \(\color{#0cbb34}{\text{End of Quote}}\) 8

TETSXPREME:

shes in 7th or 8thh im guessing

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME shes in 7th or 8thh im guessing \(\color{#0cbb34}{\text{End of Quote}}\) I am in 8th

xxDeppressionxx:

Ok so your learning about slope right?

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Ok so your learning about slope right? \(\color{#0cbb34}{\text{End of Quote}}\) yes

xxDeppressionxx:

Ok here you go–The slope will be the same for a straight line no matter which two points you pick as you know. All you need to do is to calculate the difference in the y coordinates of the 2 points and divide that by the difference of the x coordinates of the points(rise over run). That will give you the slope. Site source is (Google)

xxDeppressionxx:

Do you have any question answer choices?

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Do you have any question answer choices? \(\color{#0cbb34}{\text{End of Quote}}\) no I don't it's supposed to be a fraction

xxDeppressionxx:

The slope would be what?

carlosesgirl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx The slope would be what? \(\color{#0cbb34}{\text{End of Quote}}\) I don't get it

carlosesgirl:

@xxDeppressionxx @TETSXPREME help someone

xxDeppressionxx:

Replace all occurrences of ,, -, , - with a single +, +. Two consecutive minus signs have the same mathematical meaning as a single plus sign because 11=1-1-1=1(8,7),(1,5)(8,7),(1,5)Slope is equal to the change in y y over the change in x x, or rise over run.m= change in y change in x m= change in y change in x The change in x x is equal to the difference in x-coordinates (also called run), and the change in y y is equal to the difference in y-coordinates (also called rise).m=y2y1x2x1m=y2-y1x2-x1 Substitute in the values of x x and y y into the equation to find the slope. m=5(7)1(8)m=5-(7)1-(8)Simplify. Tap for more steps... m=2/7